OpenStudy (anonymous):

Find the inverse of (x)/(x-2)

OpenStudy (anonymous):

umm it says the answer is (2x)/(x-1) so you're wrong

OpenStudy (anonymous):

f(x) = x/(x-2) Looking at this we know that x cannot equal 2 because then the bottom would be 0. Let's change f(x) to y just to make it more simple to manipulate the equation. y= x/(x-2) Now multiply the other side to have a common denominator y(x-2) = x Multiply it out yx - 2y = x Put all the x's on one side and every thing else on the other side yx - x = 2y Keep manipulating the equation x ( y - 1 ) = 2y Now just solve for x x = 2y/ (y - 1) And for the final step just reverse the variables x for y and vice versa. y = 2x/(x-1) Do you understand what I did?

OpenStudy (anonymous):

yes and it is very nice too.

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

thanks that was really helpful

OpenStudy (anonymous):

now that you are a pro, try this \[f(x)=\frac{x}{x-2}\] \[y=\frac{x}{x-2}\] \[y(x-2)=x\] etc

OpenStudy (anonymous):

y = (x)/(x-2) x = y/(y-2) x(y-2)=y x(y-2+2)=y+2x x(y)=y+2x xy - y = 2x y(x-1)=2x y=(2x)/(x-1)